OpenStudy (danielbarriosr1):

you have $3.70 in dimes and quarters. You have 5 more quarters than dimes. how many of each type of coin do you have?

OpenStudy (anonymous):

Say the dimes be x The quarters will be x+5 The sum of their values = 370 cents 10x+25(x+5)=370 10x+25x+125=370 35x=370-125 x=245/35 x=7 (the number of dimes ) x+7= 7+5= 12 (the number of quarters)

OpenStudy (danielbarriosr1):

Thanks @sunnyshores

OpenStudy (anonymous):

np :)